# Affine Chord Sobolev Inequalities from Integral Geometry

### v1

@@ 9. Reversed inequalities
To extend Theorem 9.1 to the functional setting, we rely on the following result from [27]. It generalizes the classical result of Marshall, Olkin and Proschan [41], whose simple proof was later given by Milman and Pajor [43].
Lemma 9.2 ([27], Lemma 17).
Let $\omega:[0,\infty)\to[0,\infty)$ be decreasing with
$0<\int_{0}^{\infty}t^{\alpha-1}\omega(t)dt<\infty$
@@
$0<\int_{0}^{\infty}t^{\alpha-1}(\omega(0)-\omega(t))dt<\infty$
for $\alpha\in(-1,0)$. If $\varphi:[0,\infty)\to[0,\infty)$ is non-zero, with $\varphi(0)=0$, and such that $t\mapsto\varphi(t)$ and $t\mapsto{\varphi(t)}/{t}$ are increasing on $(0,\infty)$, then
$\zeta(\alpha)=\left\{\begin{aligned} &\left(\frac{\int_{0}^{\infty}t^{\alpha-1}\omega(\varphi(t))dt}{\int_{0}^{\infty}t^{\alpha-1}\omega(t)dt}\right)^{1/\alpha}&&\text{for}~~\alpha>0,\\ &\exp\left(\int_{0}^{\infty}\frac{\omega(\varphi(t))-\omega(t)}{t\omega(0)}dt\right)&&\text{for}~~\alpha=0,\\ &\left(\frac{\int_{0}^{\infty}t^{\alpha-1}(\omega(\varphi(t))-\omega(0))dt}{\int_{0}^{\alpha-1}(\omega(t)-\omega(0))dt}\right)^{1/\alpha}&&\text{for}~~-1<\alpha<0.\end{aligned}\right$.
is a continuous, decreasing function of $\alpha$ on $(-1,\infty)$. Moreover, $\zeta$ is constant on $(-1,\infty)$ if $\varphi(t)=\lambda t$ on $0,\infty)$ for some $\lambda>0$.
In particular, taking $\omega(t)=e^{-t}$ and $\omega(t)=(1-st)_{+}^{1/s}$ with $s>0$ yields the classical analytic continuation formulas for the Gamma and Beta functions,
@@
We note here, in the setting of Lemma 9.2, that if $\zeta(\alpha)$ is constant on $(-1,\infty)$, say $\zeta(\alpha)=\lambda$, then $\varphi(t)=\lambda t$ for all $t\geq 0$. When $\omega(t)$ is strictly decreasing, this conclusion follows from the proof of Milman and Pajor, but for completeness, we include a proof below.
@@
Let $\omega$, $\varphi$, and $\zeta$ be as in Lemma 9.2. Suppose $\omega$ is strictly decreasing and $1<\alpha<\beta<\infty$. Then $\zeta(\alpha)=\zeta(\beta)$ implies $\varphi(t)=\lambda t$ with $\lambda=\zeta(\alpha)^{-1}$.
@@
$\int_{0}^{\infty}\omega(\lambda t)t^{\alpha-1}dt=\int_{0}^{\infty}\omega(\varphi(t))t^{\alpha-1}dt.$
We set $\Psi(r)=\int_{r}^{\infty}(\omega(\lambda t)-\omega(\varphi(t)))dt$, where $\Psi(0)=\Psi(\infty)=0$. Since $\varphi(t)/t$ is increasing while $\omega$ is decreasing, we have
$\omega(\varphi(t))>\omega(\lambda t),~~\text{if}~~{\varphi(t)}/{t}<\lambda,$
@@
$\omega(\varphi(t))\leq\omega(\lambda t),~~\text{if}~~{\varphi(t)}/{t}\geq\lambda.$
This implies the sign of $\omega(\lambda r)-\omega(\varphi(r))=-\Psi^{\prime}(r)$. Therefore, $\Psi$ is first increasing and then decreasing, which yields that $\Psi(r)\geq 0$ for all $r\geq 0$. That is
$\int_{r}^{\infty}\omega(\lambda t)t^{\alpha-1}dt\geq\int_{r}^{\infty}\omega(\varphi(t))dt.$
@@
$\int_{0}^{\infty}\omega(\varphi(t))t^{\beta-1}dt$ $=(\beta-\alpha)\int_{0}^{\infty}r^{\beta-\alpha-1}\Big(\int_{r}^{\infty}\omega(\varphi(t))r^{\alpha-1}dt\Big)dr$
$\leq(\beta-\alpha)\int_{0}^{\infty}r^{\beta-\alpha-1}\Big(\int_{r}^{\infty}\omega(\lambda t)t^{\alpha-1}dt\Big)dr$
$=\int_{0}^{\infty}\omega(\lambda t)t^{\beta-1}dt=\frac{1}{\lambda^{\beta}}\int_{0}^{\infty}\omega(t)t^{\beta-1}dt,$
which implies $\zeta(\alpha)\geq\zeta(\beta)$. As we assumed, $\zeta(\alpha)=\zeta(\beta)=\lambda^{-1}$, equality holds in the inequality above, which implies that $\Psi(r)=0$ for all $r\geq 0$. Then we have
$\Psi^{\prime}(r)=\omega(\varphi(t))-\omega(\lambda t)=0.$
@@ 9.1. Results for log-concave functions
We begin this subsection by recalling a result from [2], which will serve as a main tool in the equality characterization in Theorem 3.
@@
[2, Lemma 4.2] Let $f\in L^{1}(\mathbb{R}^{n})$ be non-zero and log-concave such that $f(o)=\|f\|_{\infty}$. Let
@@
Then for every $x\in\mathbb{R}^{n}$ and $\lambda\in[0,1]$, $g(\lambda x)=g(o)^{1-\lambda}g(x)^{\lambda}$ if and only if $f(x)=\|f\|_{\infty}e^{-\|x\|_{\triangle}}$ with $\triangle$ a simplex containing the origin.
For the reader’s convenience, we restate Theorem 4 below before giving the proof.
@@
Let $f\in L^{1}(\mathbb{R}^{n})$ be non-zero and log-concave. Then
@@
for $1<\alpha<\beta<\infty$. In each inclusion, equality holds if and only if $f(x)=ae^{-\|x-x_{0}\|_{\triangle}}$ for some $a>0$ and $x_{0}\in\mathbb{R}^{n}$, where $\triangle$ a simplex containing the origin.
@@
$g(tu)=\int_{\mathbb{R}^{n}}\min\{f(x),f(x+tu)\}dx$
is an even, log-concave function and attains maximum at $t=0$.
For fixed $u\in\mathbb{S}^{n-1}$, we apply Lemma 9.2 to $\omega(t)=g(0)e^{-t}$ and $\varphi(t)=-\log(g(tu)/g(0))$. We observe that
$\zeta(\alpha)=\bigg(\frac{\int_{0}^{\infty}t^{\alpha-1}g(tu)dt}{\int_{0}^{\infty}t^{\alpha-1}g(0)e^{-t}dt}\bigg)^{1/\alpha}=\frac{\rho_{{\rm R}_{\alpha}f}(u)}{\Gamma(\alpha+1)^{1/\alpha}}$
@@
$g(0)-g(tu)$ $=\|f\|_{1}-\int_{\mathbb{R}^{n}}\min\{f(x),f(x+tu)\}dx$ (9.3)
$=\frac{1}{2}\int_{\mathbb{R}^{n}}f(x)dx+\frac{1}{2}\int_{rn}f(x+tu)dx-\int_{\mathbb{R}^{n}}\min\{f(x),f(x+tu)\}dx$
$=\frac{1}{2}\int_{\mathbb{R}^{n}}(f(x)-f(x+tu))_{+}dx+\frac{1}{2}\int_{\mathbb{R}^{n}}(f(x+tu)-f(x))_{+}dx$
$=\frac{1}{2}\int_{\mathbb{R}^{n}}|f(x)-f(x+tu)|dx,$
@@
$\zeta(\alpha)=\bigg(\frac{\int_{0}^{\infty}t^{\alpha-1}(g(0)-g(tu))dt}{\int_{0}^{\infty}t^{\alpha-1}g(0)(1-e^{-t})dt}\bigg)^{1/\alpha}=\frac{\rho_{{\rm R}_{\alpha}f}(u)}{\Gamma(\alpha+1)^{1/\alpha}}$
@@
Letting $\alpha\rightarrow-1$, Lemma 8.2 gives the right inclusion. Since $\omega(t)=e^{-t}$ is strictly decreasing, by Lemma 9.3, equality holds in (9.4) if and only if $g(tu)=g(0)e^{-\lambda t}$ for some $\lambda>0$. Then Lemma 9.4 shows that $f(x)=ae^{-\|x-x_{0}\|_{\triangle}}$ for some $a>0$ and $x_{0}\in\mathbb{R}^{n}$, where $\triangle$ is a simplex containing the origin. ∎
In fact, if we set $f(x)=e^{-\|x\|_{K}}$ for a convex body $K\subset\mathbb{R}^{n}$, then inclusion (9.2) in Theorem 9.5 reduces to the geometric case stated in Theorem 9.1. However, in this functional extension, the endpoint ${\rm R}_{\infty}f$ is $\mathbb{R}^{n}$, so the inclusion no longer extends to that limit case.
Furthermore, combining Lemma 9.2 with (9.4), we obtain an explicit expression for $\zeta(0)$ by the continuity of $\zeta(\alpha)$, where
$\log\zeta(0)=-\gamma+\int_{0}^{\infty}\frac{1}{t}\Big(\frac{1}{\|f\|_{1}}\int_{\mathbb{R}^{n}}\min\{f(x),f(x+tu)\}dx-e^{-t}\Big)dt.$ (9.5)
When $f=\operatorname{1}_{K}$, the quantity $\log\zeta(0)=\log\rho_{{\rm R}_{0}K}(u)$. The next result confirms that the formula (9.5) is consistent with the definition of ${\rm R}_{0}f$ (8.2), namely $\rho_{{\rm R}_{0}f}(u)=\zeta(0)$ with $u\in\mathbb{S}^{n-1}$ fixed.
Proposition 9.6.
For fixed $u\in\mathbb{S}^{n-1}$ and non-zero, log-concave $f\in L^{1}(\mathbb{R}^{n})$,
$\rho_{{\rm R}_{0}f}(u)=\zeta(0).$
Proof.
By (9.5),
$\log\zeta(0)=$ $-\gamma+\int_{0}^{\infty}\frac{1}{t}\Big(\frac{1}{\|f\|_{1}}\int_{\mathbb{R}^{n}}\min\{f(x),f(x+tu)\}dx-e^{-t}\Big)dt$
$=-\gamma+\int_{0}^{\infty}\frac{1}{t}\Big(\int_{0}^{\|f\|_{\infty}}|\frac{\{f\geq r\}\cap(\{f\geq r\}+tu)|}{|\{f\geq r\}|}d\mu_{f}(r)-e^{-t}\Big)dt$
$=\int_{0}^{\|f\|_{\infty}}\Big(-\gamma+\int_{0}^{\infty}\frac{1}{t}\Big(\frac{|\{f\geq r\}\cap(\{f\geq r\}+tu)|}{|\{f\geq r\}|}-e^{-t}\Big)dt\Big)d\mu_{f}(r).$
Note that
$\log\rho_{{\rm R}_{0}\{f\geq r\}}(u)$ $=-\gamma+\int_{0}^{\infty}\frac{1}{t}\Big(\frac{1}{|\{f\geq r\}|}\int_{\mathbb{R}^{n}}\min\{\operatorname{1}_{\{f\geq r\}}(x),\operatorname{1}_{\{f\geq r\}}(x+tu)\}dx-e^{-t}\Big)dt$
$=-\gamma+\int_{0}^{\infty}\frac{1}{t}\Big(\frac{|\{f\geq r\}\cap(\{f\geq r\}+tu)|}{|\{f\geq r\}|}-e^{-t}\Big)dt,$
where the first equation is from (9.5). Therefore, we have
$\log\zeta(0)=\int_{0}^{\|f\|_{\infty}}\log\rho_{{\rm R}_{0}\{f\geq r\}}(u)d\mu_{f}(r)=\log\rho_{{\rm R}_{0}f}(u),$
which concludes the proof. ∎
@@ 9.2. Results for $s$-concave functions
In this subsection, we focus on the case of $s$-concave functions. The following lemma is a direct corollary from the Borell-Brascamp-Lieb inequality, which is also mentioned in [15, Theorem 11.2].
@@
For $s$-concave $f\in L^{1}(\mathbb{R}^{n})$ and fixed $u\in\mathbb{S}^{n-1}$,
$t\mapsto g(tu)=\int_{\mathbb{R}^{n}}\min\{f(x),f(x+tu))\}dx$
is a $s/(ns+1)$-concave function on $\mathbb{R}$.
@@
$N_{f}(x,t)=\min\{f(x),f(x+tu)\}.$
First, we prove that $N_{f}(x,t)$ is $s$-concave on $\mathbb{R}^{n}\times\mathbb{R}$. For $x_{0},x_{1}\in\mathbb{R}^{n}$, $t_{0},t_{1}\in\mathbb{R}$ and $\lambda\in[0,1]$, we denote $x_{\lambda}=(1-\lambda)x_{0}+\lambda x_{1}$ and $t_{\lambda}=(1-\lambda)t_{0}+\lambda t_{1}$. Note that
$N_{f}(x_{\lambda},t_{\lambda})^{s}=\min\{f(x_{\lambda})^{s},f(x_{\lambda}+t_{\lambda}u)^{s}\}.$
@@
$N_{f}(x_{\lambda},t_{\lambda})^{s}$ $\geq\min\{(1-\lambda)f(x_{0})^{s}+\lambda f(x_{1})^{s},(1-\lambda)f(x_{0}+t_{0}u)^{s}+\lambda f(x_{1}+t_{1}u)^{s}\}$
$\geq(1-\lambda)\min\{f(x_{0})^{s},f(x_{0}+t_{0}u)^{s}\}+\lambda\min\{f(x_{1})^{s},f(x_{1}+t_{1}u)^{s}\}$
$=(1-\lambda)N_{f}(x_{0},t_{0})^{s}+\lambda N_{f}(x_{1},t_{1})^{s},$
@@
For $\lambda\in[0,1]$, and fixed $t_{0},t_{1}\in\mathbb{R}$, we set
$F(x)=N_{f}(x,t_{0}),~~G(x)=N_{f}(x,t_{1}),~~H(x)=N_{f}(x,t_{\lambda}),$
where $t_{\lambda}=(1-\lambda)t_{0}+\lambda t_{1}$. By the $s$-concavity of $F$, we have
$H((1-\lambda)x+\lambda y)\geq\mathcal{M}_{\lambda}^{s}(F(x),G(y)).$
The Borell-Brascamp-Lieb inequality (2.3) then implies that
@@
that is, the $\frac{s}{ns+1}$-concavity of $g(tu)$ in $t$. ∎
We begin the proof of Theorem 5 by stating its inclusion result as the following theorem; the equality case will be addressed separately.
@@
Let $f\in L^{1}(\mathbb{R}^{n})$ be $s$-concave with compact support. If $1<\beta<\alpha$, then
@@
By Lemma 9.7, the function $g(tu)=\int_{\mathbb{R}^{n}}\min\{f(x),f(x+tu)\}dx$ is $s^{\prime}$-concave with $s^{\prime}=s/(ns+1)$. We apply Lemma 9.2 with $\omega(t)=g(0)(1-s^{\prime}t)_{+}^{1/s^{\prime}}$ and $\varphi(t)=\frac{1}{s^{\prime}}(1-(g(tu)/g(0))^{s^{\prime}})$.
@@
$\zeta(\alpha)=\bigg(\frac{\int_{0}^{\infty}t^{\alpha-1}g(tu)dt}{\int_{0}^{\infty}g(0)t^{\alpha-1}(1-s^{\prime}t)_{+}^{1/s^{\prime}}dt}\bigg)^{1/\alpha}=\frac{s^{\prime}\rho_{{\rm R}_{\alpha}f}(u)}{(\alpha B(\alpha,1+\frac{1}{s^{\prime}}))^{1/\alpha}}$
@@
$\zeta(\alpha)=\bigg(\frac{\int_{0}^{\infty}t^{\alpha-1}(g(0)-g(tu))dt}{\int_{0}^{\infty}g(0)t^{\alpha-1}((1-s^{\prime}t)_{+}^{1/s^{\prime}}-1)dt}\bigg)^{1/\alpha}=\frac{s^{\prime}\rho_{{\rm R}_{\alpha}f}(u)}{(\alpha B(\alpha,1+\frac{1}{s^{\prime}}))^{1/\alpha}}$
@@
$\Big(\alpha B(\alpha,1+\frac{1}{s^{\prime}})\Big)^{-1/\alpha}=\Big(nB(\alpha+1,\frac{1}{s^{\prime}})\Big)^{-1/\alpha}=c_{n,\alpha}(s).$
@@
Since Lemma 8.3 ensures that ${\rm R}_{\beta}f\to{\rm R}_{\infty}f$ and $c_{n,\beta}(s)\to 1$ as $\beta\rightarrow\infty$, the left inclusion follows directly from (9.8).

### v2

@@ 6. Reverse inclusion for radial mean bodies
When $\beta=n$, the left-most inclusion implies the Rogers-Shephard inequality:
$|{\rm D}K|\leq\binom{2n}{n}|K|.$
When $\alpha=n$, the right-most inclusion yields Zhang’s inequality:
where $\triangle$ denotes a simplex in $\mathbb{R}^{n}$.
To extend Theorem 6.1 to the functional setting, we rely on the following result from [26]. It generalizes the classical result of Marshall, Olkin and Proschan [38], whose proof was later simplified by Milman and Pajor [40].
Lemma 6.2.
[26, Lemma 17] Let $\omega:[0,\infty)\to[0,\infty)$ be decreasing with
$0<\int_{0}^{\infty}r^{\alpha-1}\omega(r)dr<\infty$
@@
$0<\int_{0}^{\infty}r^{\alpha-1}(\omega(0)-\omega(r))dr<\infty$
for $\alpha\in(-1,0)$. If $\varphi:[0,\infty)\to[0,\infty)$ is non-zero, with $\varphi(0)=0$, and such that $r\mapsto\varphi(r)$ and $r\mapsto{\varphi(r)}/{r}$ are increasing on $(0,\infty)$, then
$\zeta(\alpha)=\left\{\begin{aligned} &\left(\frac{\int_{0}^{\infty}r^{\alpha-1}\omega(\varphi(r))dr}{\int_{0}^{\infty}r^{\alpha-1}\omega(r)dr}\right)^{1/\alpha}&&\text{for}~~\alpha>0,\\ &\exp\left(\int_{0}^{\infty}\frac{\omega(\varphi(r))-\omega(r)}{r\omega(0)}dr\right)&&\text{for}~~\alpha=0,\\ &\left(\frac{\int_{0}^{\infty}r^{\alpha-1}(\omega(\varphi(r))-\omega(0))dr}{\int_{0}^{\alpha-1}(\omega(r)-\omega(0))dr}\right)^{1/\alpha}&&\text{for}~~-1<\alpha<0,\end{aligned}\right$.
is a continuous, decreasing function of $\alpha$ on $(-1,\infty)$. Moreover, $\zeta$ is constant on $(-1,\infty)$ if $\varphi(r)=\lambda r$ on $0,\infty)$ for some $\lambda>0$.
In particular, taking $\omega(r)=e^{-r}$ and $\omega(r)=(1-sr)_{+}^{1/s}$ with $s>0$ yields the classical analytic continuation formulas for the Gamma and Beta functions,
@@
Under the assumptions of Lemma 6.2 with strictly decreasing $\omega(r)$, if $\zeta(\alpha)$ is constant on $(-1,\infty)$, say $\zeta(\alpha)=\lambda$, then $\varphi(r)=\lambda r$ for all $r\geq 0$. The result follows directly from Milman and Pajor [40], but we provide a proof for completeness.
@@
Let $\omega$, $\varphi$, and $\zeta$ be as in Lemma 6.2. Suppose $\omega$ is strictly decreasing and $1<\alpha<\beta<\infty$. Then $\zeta(\alpha)=\zeta(\beta)$ implies $\varphi(r)=\lambda r$ with $\lambda=\zeta(\alpha)^{-1}$.
@@
$\int_{0}^{\infty}\omega(\lambda r)r^{\alpha-1}dr=\int_{0}^{\infty}\omega(\varphi(r))r^{\alpha-1}dr.$
We set $\Psi(t)=\int_{t}^{\infty}(\omega(\lambda r)-\omega(\varphi(r)))dr$, where $\Psi(0)=\Psi(\infty)=0$. Since $\varphi(r)/r$ is increasing while $\omega$ is decreasing, we have
$\omega(\varphi(r))>\omega(\lambda r),~~\text{if}~~{\varphi(r)}/{r}<\lambda,$
@@
$\omega(\varphi(r))\leq\omega(\lambda r),~~\text{if}~~{\varphi(r)}/{r}\geq\lambda.$
This implies the sign of $\omega(\lambda t)-\omega(\varphi(t))=-\Psi^{\prime}(t)$. Therefore, $\Psi$ is first increasing and then decreasing, which yields that $\Psi(t)\geq 0$ for all $t\geq 0$. That is
$\int_{t}^{\infty}\omega(\lambda r)r^{\alpha-1}dr\geq\int_{t}^{\infty}\omega(\varphi(r))r^{\alpha-1}dr.$
@@
$\int_{0}^{\infty}\omega(\varphi(r))r^{\beta-1}dr$ $=(\beta-\alpha)\int_{0}^{\infty}t^{\beta-\alpha-1}\Big(\int_{t}^{\infty}\omega(\varphi(r))r^{\alpha-1}dr\Big)dt$
$\leq(\beta-\alpha)\int_{0}^{\infty}t^{\beta-\alpha-1}\Big(\int_{t}^{\infty}\omega(\lambda r)r^{\alpha-1}dr\Big)dt$
$=\int_{0}^{\infty}\omega(\lambda r)r^{\beta-1}dr=\frac{1}{\lambda^{\beta}}\int_{0}^{\infty}\omega(r)r^{\beta-1}dr,$
which implies $\zeta(\alpha)\geq\zeta(\beta)$. If $\zeta(\alpha)=\zeta(\beta)=\lambda^{-1}$, equality holds in the inequality above, which yields that $\Psi(t)=0$ for all $t\geq 0$. Then we have
$\Psi^{\prime}(t)=\omega(\varphi(t))-\omega(\lambda t)=0.$
@@ 6.1. Results for log-concave functions
In this subsection, for the sake of completeness, we provide a proof of (1.13), which extends (1.12) to the class of log-concave functions and was recently established in [28, Theorem 1.6]. The following result from [3] serves as a main tool in the equality characterization of (1.13).
@@
[3, Lemma 4.2] Let $f\in L^{1}(\mathbb{R}^{n})$ be non-zero and log-concave such that $f(o)=\|f\|_{\infty}$. Let
@@
Then for every $x\in\mathbb{R}^{n}$ and $\lambda\in(0,1)$, $g(\lambda x)=g(o)^{1-\lambda}g(x)^{\lambda}$ if and only if $f(x)=\|f\|_{\infty}e^{-\|x\|_{\triangle}}$ with $\triangle$ a simplex containing the origin.
For the reader’s convenience, we restate (1.13) below and give a proof.
@@
Let $f\in L^{1}(\mathbb{R}^{n})$ be a non-zero, log-concave function. Then
@@
for $1<\alpha<\beta<\infty$. In each inclusion, there is equality if and only if $f(x)=ae^{-\|x-x_{0}\|_{\triangle}}$ where $a=\|f\|_{\infty}$ and $\triangle$ is a simplex containing $x_{0}$.
@@
$g(ru)=\int_{\mathbb{R}^{n}}\min\{f(x),f(x+ru)\}dx$
is an even, log-concave function and attains maximum at $r=0$.
For fixed $u\in\mathbb{S}^{n-1}$, we apply Lemma 6.2 to $\omega(r)=g(o)e^{-r}$ and $\varphi(r)=-\log(g(ru)/g(o))$. We observe that
$\zeta(\alpha)=\bigg(\frac{\int_{0}^{\infty}r^{\alpha-1}g(ru)dr}{\int_{0}^{\infty}r^{\alpha-1}g(o)e^{-r}dr}\bigg)^{1/\alpha}=\frac{\rho_{{\rm R}_{\alpha}f}(u)}{\Gamma(\alpha+1)^{1/\alpha}}$
@@
$g(o)-g(ru)$ $=\|f\|_{1}-\int_{\mathbb{R}^{n}}\min\{f(x),f(x+ru)\}dx$ (6.3)
$=\frac{1}{2}\int_{\mathbb{R}^{n}}f(x)dx+\frac{1}{2}\int_{\mathbb{R}^{n}}f(x+ru)dx-\int_{\mathbb{R}^{n}}\min\{f(x),f(x+ru)\}dx$
$=\frac{1}{2}\int_{\mathbb{R}^{n}}(f(x)-f(x+ru))_{+}dx+\frac{1}{2}\int_{\mathbb{R}^{n}}(f(x+ru)-f(x))_{+}dx$
$=\frac{1}{2}\int_{\mathbb{R}^{n}}|f(x)-f(x+ru)|dx,$
@@
$\zeta(\alpha)=\bigg(\frac{\int_{0}^{\infty}r^{\alpha-1}(g(o)-g(ru))dr}{\int_{0}^{\infty}r^{\alpha-1}g(o)(1-e^{-r})dr}\bigg)^{1/\alpha}=\frac{\rho_{{\rm R}_{\alpha}f}(u)}{\Gamma(\alpha+1)^{1/\alpha}}$
@@
Letting $\alpha\rightarrow-1$, Lemma 5.6 gives the right inclusion. Since $\omega(r)=e^{-r}$ is strictly decreasing, by Lemma 6.3, equality holds in (6.4) if and only if $g(ru)=g(o)e^{-\lambda r}$ for some $\lambda>0$. Then Lemma 6.4 shows that $f(x)=ae^{-\|x-x_{0}\|_{\triangle}}$ for $a=\|f\|_{\infty}$ and $x_{0}\in\triangle$, where $\triangle$ is a simplex containing the origin. ∎
If we set $f(x)=e^{-\|x\|_{K}}$ for a convex body $K\subset\mathbb{R}^{n}$, then inclusion (6.2) in Theorem 6.5 reduces to the geometric case stated in Theorem 6.1. However, in this functional extension the endpoint ${\rm R}_{\infty}f$ is $\mathbb{R}^{n}$, so the inclusion no longer extends to that limit.
This observation motivates the consideration of the embedding $K\mapsto\operatorname{1}_{K}$ in studying functional extensions of Theorem 6.1. The following counterexample shows, however, that this extension fails for general log-concave functions, which naturally leads to the study of $s$-concave functions in the next subsection.
Example 6.6.
Let $n=1$ and $f(x)=2x\operatorname{1}_{[0,1]}(x)$. It is clear that $\log f(x)=\log x+\log 2$ for $x\in[0,1]$, which is concave on ${\rm supp}(f)$. Then we have
$g(r)=(1-|r|)_{+}^{2}.$
In this case, we consider $\beta=1$ and $\alpha=\frac{1}{2}$ and obtain that
$c_{1,1}{\rm R}_{1}f=\Big[-\frac{2}{3},\frac{2}{3}\Big],\quad c_{1,\frac{1}{2}}{\rm R}_{\frac{1}{2}}f=\Big[-\frac{16}{25},\frac{16}{{25}}\Big],$
@@ 6.2. Results for $s$-concave functions
In this subsection, we focus on the case of $s$-concave functions with $s>0$. We continue to use the notation
$g(ru)=\int_{\mathbb{R}^{n}}\min\{f(x),f(x+ru)\}dx.$
The following lemma is a direct corollary from the Borell-Brascamp-Lieb inequality, which is also mentioned in [17, Theorem 11.2].
@@
Let $s>0$ and $u\in\mathbb{S}^{n-1}$ be fixed. For an $s$-concave, integrable function $f$, the function $r\mapsto g(ru)$ is $s/(ns+1)$-concave on $\mathbb{R}$.
@@
$N_{f}(x,r)=\min\{f(x),f(x+ru)\}.$
First, we prove that $N_{f}(x,r)$ is $s$-concave on $\mathbb{R}^{n}\times\mathbb{R}$. For $x_{0},x_{1}\in\mathbb{R}^{n}$, $r_{0},r_{1}\in\mathbb{R}$ and $\lambda\in[0,1]$, we denote $x_{\lambda}=(1-\lambda)x_{0}+\lambda x_{1}$ and $r_{\lambda}=(1-\lambda)r_{0}+\lambda r_{1}$. Note that
$N_{f}(x_{\lambda},r_{\lambda})^{s}=\min\{f(x_{\lambda})^{s},f(x_{\lambda}+r_{\lambda}u)^{s}\}.$
@@
$N_{f}(x_{\lambda},r_{\lambda})^{s}$ $\geq\min\{(1-\lambda)f(x_{0})^{s}+\lambda f(x_{1})^{s},(1-\lambda)f(x_{0}+r_{0}u)^{s}+\lambda f(x_{1}+r_{1}u)^{s}\}$
$\geq(1-\lambda)\min\{f(x_{0})^{s},f(x_{0}+r_{0}u)^{s}\}+\lambda\min\{f(x_{1})^{s},f(x_{1}+r_{1}u)^{s}\}$
$=(1-\lambda)N_{f}(x_{0},r_{0})^{s}+\lambda N_{f}(x_{1},r_{1})^{s},$
@@
For $\lambda\in[0,1]$, and fixed $r_{0},r_{1}\in\mathbb{R}$, we set
$F(x)=N_{f}(x,r_{0}),~~G(x)=N_{f}(x,r_{1}),~~H(x)=N_{f}(x,r_{\lambda}),$
where $r_{\lambda}=(1-\lambda)r_{0}+\lambda r_{1}$. By the $s$-concavity of $N_{f}$, we have
$H((1-\lambda)x+\lambda y)\geq\mathcal{M}_{\lambda}^{s}(F(x),G(y)),$
where $\mathcal{M}_{\lambda}^{s}$ denotes the $s$-mean (see Section 2.4). The Borell-Brascamp-Lieb inequality (2.4) then implies that
@@
that is, the $\frac{s}{ns+1}$-concavity of $g(ru)$ in $r$. ∎
We begin the proof of Theorem 3 by stating the inclusion result as the following theorem, the equality case will be addressed separately.
@@
Let $f\in L^{1}(\mathbb{R}^{n})$ be a non-zero $s$-concave function with compact support. If $1<\alpha<\beta$, then
@@
By Lemma 6.7, the function $g(ru)=\int_{\mathbb{R}^{n}}\min\{f(x),f(x+ru)\}dx$ is $s^{\prime}$-concave with $s^{\prime}=s/(ns+1)$. We apply Lemma 6.2 with $\omega(r)=g(0)(1-s^{\prime}r)_{+}^{1/s^{\prime}}$ and $\varphi(r)=\frac{1}{s^{\prime}}(1-(g(ru)/g(0))^{s^{\prime}})$.
@@
$\zeta(\alpha)=\bigg(\frac{\int_{0}^{\infty}r^{\alpha-1}g(ru)dr}{\int_{0}^{\infty}g(0)r^{\alpha-1}(1-s^{\prime}r)_{+}^{1/s^{\prime}}dr}\bigg)^{1/\alpha}=\frac{s^{\prime}\rho_{{\rm R}_{\alpha}f}(u)}{(\alpha B(\alpha,1+\frac{1}{s^{\prime}}))^{1/\alpha}}$
@@
$\zeta(\alpha)=\bigg(\frac{\int_{0}^{\infty}r^{\alpha-1}(g(0)-g(ru))dr}{\int_{0}^{\infty}g(0)r^{\alpha-1}((1-s^{\prime}r)_{+}^{1/s^{\prime}}-1)dr}\bigg)^{1/\alpha}=\frac{s^{\prime}\rho_{{\rm R}_{\alpha}f}(u)}{(\alpha B(\alpha,1+\frac{1}{s^{\prime}}))^{1/\alpha}}$
@@
$\Big(\alpha B(\alpha,1+\frac{1}{s^{\prime}})\Big)^{-1/\alpha}=\Big(\frac{1}{s^{\prime}}B(\alpha+1,\frac{1}{s^{\prime}})\Big)^{-1/\alpha}=c_{n,\alpha}(s).$
@@
Since Lemma 5.7 ensures that ${\rm R}_{\beta}f\to{\rm R}_{\infty}f$ and $c_{n,\beta}(s)\to 1$ as $\beta\rightarrow\infty$, the left-most inclusion follows directly from (6.7) by letting $\beta\rightarrow\infty$.